Given a student has scored

M marks in Maths, P marks in Physics and C marks in Chemistry. Write a program to check if a student is eligible for admission in a professional course based on the following criteria:Sum of marks in any two subjects >= 100 and M + P + C >= 180.

Input

The first line is an integer

M, the second line is an integer P, and the third line is an integer C.Output

The output should be a single line containing

True if it satisfies the given conditions, False in all other cases.Explanation


Given

M = 82, P = 55, C = 45.


M + P >= 100,(137 >= 100)

M + C >= 100,(127 >= 100)

and, M + P + C >= 180 (82 + 55 + 45 >= 180)

 


So, the output should be

True.

Sample Input 1

82

55

45

Sample Output 1

True

Sample Input 2

71

30

70

Sample Output 2

False


1
Expert's answer
2021-09-06T23:58:13-0400
M=int(input("Maths: "))
P=int(input("Physics: "))
C=int(input("Chemistry: "))
if ((M+P)>=100 or (M+C)>=100 or (P+C)>=100) and (M+P+C)>=180:
    print("True")
else:
    print("False")